OpenStudy (anonymous):

(This is a matrix equation and I'm so confused)Find the variables: [-1 -w^2] = [ 3k 1] [12f 0 ] [ 24 12]

OpenStudy (anonymous):

-1=3k 12f=24 -w^2=1 solve this

OpenStudy (anonymous):

k = -1/3 f = 2 w = i Is that right? Also, how do you get here, I thought you were supposed to do the whole thing across like -1 - w^2 = 3k + 1

OpenStudy (anonymous):

two matrixes are equal if they have same elements. That's it

OpenStudy (anonymous):

Oh! I get it now, thanks a ton

OpenStudy (anonymous):

yw